# High-dimensional stochastic finite volumes using the tensor train format

### v1

@@ 5.2.2 Parameter study
In summary, we showed that for the Sod problem, a low rank representation is sufficient in most places, but higher rank are necessary in some local parts. This example illustrates how we can use the mixed format to gain more precise information on the problem at hand and tune the parameters to improve the solution. For simplicity, we chose here to keep the same rank for all spatial cells, but the algorithm could easily be changed to have a higher rank only in some cells.
6 Conclusion
In this paper, we have introduced a new method to use the stochastic finite volume method with many stochastic parameters. The numerical experiments show the feasibility of the approach. They also underline the importance of correctly choosing the TT-rank for the tensor approximation: there is a trade-off between the accuracy and the need to keep computational cost tractable.
@@ 6 Conclusion
  • The choice of correct algorithm parameters – such as the TT-rank or the tolerance for the cross approximation – is very likely to be problem-dependent. Hence, the algorithm should be tested on other representative problems and for various probability densities.
  • The convergence order of deterministic methods is altered by the introduction of random parameters. This aspect should be further investigated to provide convergence estimates for this recent class of algorithms.
  • The method proposed in this paper keeps each spatial cell separate, which should allow some analysis. In particular, it would be interesting to look for correlations between the features in the physical space and the TT-ranks in stochastic space. From the Sod example, we assume that imposing higher ranks only on few cells should improve the result while keeping computational cost low. Further investigations are required to validate this hypothesis.
A comparison with the fully compressed approach [7] would also be of interest. One could for example compare the number of parameters needed by both approaches to obtain the same accuracy, when using the same finite volume method and the same propagation in time.
@@
The proposed algorithm is a proof of concept and can easily be combined with various high-performance computing techniques. A mesh adaptation in the physical space could take into account the TT-rank. For example, the method introduced in [21, 22, 17] for mesh adaptation in physical and stochastic space could be adapted to our setting. Moreover, the algorithm is easily parallelizable, since the numerical flux is computed independently on each spatial cell.
Aknowledgements
The authors thank Adrian Kolb for the Monte-Carlo results. The work has been supported by the DFG (German Research Foundation) through 320021702/GRK2326: Energy, Entropy, and Dissipative Dynamics (EDDy) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - SPP 2410 Hyperbolic Balance Laws in Fluid Mechanics: Complexity, Scales, Randomness (CoScaRa) within the Projects 525842915, 52584291 (Numerische Verfahren für gekoppelte Mehrskalenprobleme) and 525853336 (Zufällige kompressible Euler Gleichungen: Numerik und ihre Analysis). The authors thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for the financial support through 333849990/GRK2379 (IRTG Hierarchical and Hybrid Approaches in Modern Inverse Problems). Also, support received funding from the European Union’s Horizon Europe research and innovation program under the Marie Sklodowska-Curie Doctoral Network Datahyking (Grant No. 101072546).
Appendix A Algorithms.
@@
We describe the method in pseudocode. The general algorithm is presented in Algorithm 1 and the flux computation is presented in Algorithm 2. Every time cross interpolation is used, it is made visible in the pseudocode with the syntax ’$\textsc{cross}(f(T1,T2,\dots))$’. Here, ’f’ is the non-polynomial function and ’T1, T2, …’ are its arguments in the tensor train format.
Remark 1
@@
In Algorithm 1, the maximum of $|f^{\prime}(\bar{U}^{0})|$ can be computed directly in the tensor train format. In tntorch – the tensor train library used for this work – the minimum of a tensor train is computed by minimizing the objective function $x\mapsto\tan(\pi/2-x)$, where the objective function is evaluated with cross approximation.
Algorithm 1 Algorithm for a forward Euler scheme in time
@@
     $U^{0}_{i_{q}}\leftarrow\textsc{cross}(u_{0}(x_{i_{q}},\xi_{1},\dots,\xi_{m}))$triangleright Evaluate $u_{0}$ on the quadrature grid with cross approximation
@@
$\bar{U}^{0}\leftarrow\textsc{Average}(U^{0})$triangleright Compute the cell average of the initial condition $\bar{U}^{0}$
Compute $\Delta t<\frac{1}{2}\dfrac{\Delta x}{\max|f^{\prime}(\bar{U}^{0})|}$.
@@
         $U^{n\pm}_{i\pm 1/2}\leftarrow\textsc{Interpolate}(\bar{U}_{i}^{n})$triangleright Interpolate values at interfaces, see Algorithm 3
         $H_{i\pm 1/2}\leftarrow\textsc{NumFlux}(U^{\pm}_{i\pm 1/2})$triangleright Compute the numerical flux at interfaces, see Algorithm 2
         $\bar{U}_{i}^{n+1}\leftarrow\bar{U}_{i}^{n}-\frac{\Delta t}{\Delta x}(H_{i+1/2}-H_{i-1/2})$triangleright Update the solution
@@
Algorithm 2 Numerical flux computation for a MUSCL-type scheme
$a^{+}\leftarrow\textsc{cross}(\max(|f^{\prime}(U_{i+1/2}^{+})|,|f^{\prime}(U_{i+1/2}^{-})|))$
$H_{i+1/2}\leftarrow\dfrac{f(U_{i+1/2}^{+})+f(U_{i+1/2}^{-})}{2\Delta x}-\dfrac{a^{+}(U_{i+1/2}^{+}-U_{i+1/2}^{-})}{2\Delta x}$
return $H_{i+1/2}$
Algorithm 3 Interpolation
@@
     $(U_{x})_{i}^{n}\leftarrow\textsc{cross}(\text{minmod}(U_{i}-U_{i-1},U_{i+1}-U_{i}))$
@@
return $U^{n+}_{i},U^{n-}_{i}$
@@ Appendix B Expectation and variance for the exact solution
$\hat{\bm{\xi}}=(\xi_{1},\dots,\xi_{m-1}),\quad\hat{\bf v}_{L}=(v_{L,1},\dots,v_{L,m-1}),\quad\hat{\bf v}_{R}=(v_{R,1},\dots,v_{R,m-1}),\quad\hat{\bf v}=(\hat{\bf v}_{L}+\hat{\bf v}_{R})/2.$ (46)
@@
$a(t,x,\hat{\bm{\xi}})=\begin{cases}1,&c(\hat{\bm{\xi}},x,t)>1,\\ c(t,x,\hat{\bm{\xi}}),&0<c(t,x,\hat{\bm{\xi}})<1,\\ 0,&c(t,x,\hat{\bm{\xi}})<0\end{cases}=\min\left(1,\max\left(0,c(t,x,\hat{\bm{\xi}})\right)\right).$ (47)
@@
$\mathbb{E}[u]=\int_{[0,1]^{m-1}}\int_{\xi_{m}=0}^{a(\hat{\bm{\xi}})}u_{R}({\bm{\xi}})p_{1}(\xi_{1})\dots p_{m}(\xi_{m})\;\text{d}\xi_{m}\;\text{d}\hat{\bm{\xi}}+\int_{[0,1]^{m-1}}\int_{\xi_{m}=a(\hat{\bm{\xi}})}^{1}u_{L}({\bm{\xi}})p_{1}(\xi_{1})\dots p_{m}(\xi_{m})\;\text{d}\xi_{m}\;\text{d}\hat{\bm{\xi}}.$ (48)

### v2

@@ 5.2.3 Comparison with the full-TT format
For the Sod problem, a low rank representation is sufficient in most places, but higher rank are necessary in some local parts. This example illustrates how we can use the hybrid format to gain more precise information on the problem.
5.3 Shu-Osher problem
The last experiment is the Shu-Osher problem for the Euler system (45)-(46). In this test case, a shock wave interacts with a sine wave density. The solution includes a shock wave and small-scale variations. Two stochastic variables are added in the initial data to extend the dimension. For $\xi_{1}=\xi_{2}=0$, we recover the the classical definition of the problem. The initial condition for the random problem in primitive variables are given by
$\begin{pmatrix}\rho(0,x,{\bm{\xi}}),\\ u(0,x,{\bm{\xi}}),\\ p(0,x,{\bm{\xi}})\end{pmatrix}=\begin{cases}\begin{pmatrix}3.857143+0.1\xi_{1}\\ 2.629369+0.1\xi_{1}\\ 10.33333+\xi_{1}\end{pmatrix},&0<x<\frac{1}{8}\\[20.0pt] \begin{pmatrix}1+0.2(1+\xi_{2})\sin(16\pi x)\\ 0.0\\ 1.+0.1\xi_{2}\end{pmatrix},&\frac{1}{8}<x<1\end{cases}.$ (50)
The simulation was run until time $t=0.13$. The following parameters are used for the discretization: $N_{x}=1600,N_{\xi}=20,{r_{\text{max}}}=5,\varepsilon=0.01$. We also run tests with a smaller maximal rank $r_{\text{max}}$} and a larger $\varepsilon$, but the solution in that case yields negative densities. Therefore, those results are omitted. The relatively high rank indicates that a strong compression of the solution is not necessarily expected for this problem.
Figure 14 shows the expectation and the standard deviation $\sigma$ for $\rho,u,p$. We see that the standard deviation is large near the shock, in particular in the results for the density. Figure 15 shows the maximal ranks in each spatial cell at the final time for $\rho,u,p$. Interestingly enough, most cells have a very low rank: $r\leq 2$. The largest ranks are, as in the Sod case, concentrated near the shock. This is also expected due to the low regularity of the solution at that point. The maximal rank at intermediate times are plotted in Appendix C. As observed in the final time, the high ranks follow the shock. For the velocity, the region to the right of the shock has a larger rank than to the left of the shock. Overall, this test case has the same general behavior as the Sod problem: ranks are kept very low for most spatial cells, and the rank increases near the shock. However, the maximal rank is still small compared to the dimension of the problem and this show the usability of the hybrid format even for more complex test-cases.
Figure 14: Expectation and standard deviation for $\rho,u,p$ at final time. The results were obtained with $N_{x}=1600,N_{\xi}=20,{r_{\text{max}}}=5,\varepsilon=0.01$.
Figure 15: Maximal rank in each spatial cell for $\rho,u,p$ at final time. The results were obtained with $N_{x}=1600,N_{\xi}=20,{r_{\text{max}}}=5,\varepsilon=0.01$.
6 Conclusion and Outlook
In this paper, we have introduced a new method to use the stochastic finite volume method with many stochastic parameters. The numerical experiments show the feasibility of the approach. A comparison with the full-TT format shows that both methods yield qualitatively similar results. In the current implementation, the full-TT format is more efficient, both in time and number of coefficients. Our current implementation, which uses regular grids and no parallelization, could be easily improved. The hybrid method also has the unique feature that the rank varies in each spatial cell, which allows more adaption to the correlation between the spatial and stochastic variables. It furthermore simply generalizes to arbitrary finite–volume schemes.
@@ 6 Conclusion and Outlook
  • The choice of appropriate algorithm parameters – such as the TT-rank or the tolerance for TT approximations – is very likely to be problem-dependent. Hence, the algorithm should be tested on other representative problems and for various probability densities.
  • Convergence estimates combining the SFV and the TT approximation are yet to be developed. Our results indicate that the tolerance for the TT approximation has some impact on the convergence rate. Their theoretical estimates is also an ongoing work.
  • From the Sod problem and the Shu-Osher problem, it appears that the ranks are generally kept very low, and that larger ranks are required near the shock. The correlation between spatial features and stochastic feature is an interesting topic for future research.
@@
The proposed algorithm is a proof of concept and it may easily be combined with various high-performance computing techniques. The algorithm is easily parallelizable, since the numerical flux is computed independently on each spatial cell. Moreover, the observations from the Sod problem and the Shu-Osher problem indicate a new direction for mesh adaptation methods. The area where the rank increase is also the region where mesh refinement would be the most beneficial: a large rank could be an indicator as to where refinement is needed. For example, the method introduced in [21, 22, 17] for mesh adaptation in physical and stochastic space may be adapted to our setting.
Appendix A Algorithms
We describe the method in pseudocode. The general algorithm for the hybrid format is presented in Algorithm 1 and the reconstruction is presented in Algorithm 2. The general algorithm and reconstruction for the full-TT format are presented in Algorithm 3 and Algorithm 4, respectively. The CFL number is denoted by $\alpha$.
In the full-TT format, the left and right numerical fluxes are computed as follows: the numerical flux at each interface is first computed using the global operations, then the left-shifted flux is computed by applying the matrix $\bf L$} to the spatial core of the flux. Those operations are described in Algorithm 5.
@@ Appendix A Algorithms
Every time cross interpolation is used, it is made visible in the pseudocode with the syntax $\text{cross}\,(f(T1,T2,\dots))$’. Here, ’f’ is the non-polynomial function and ’T1, T2, …’ are its arguments in the tensor train format.
Remark 1.
@@
In the Algorithms 1 and 3, the maximum of $|f^{\prime}(\bar{U}^{0})|$ can be computed directly in the tensor train format. In tntorch – the tensor train library used for this work – the minimum of a tensor train is computed by minimizing the objective function $x\mapsto\tan(\pi/2-x)$, where the objective function is evaluated with cross approximation.
Algorithm 1 MUSCL scheme and forward Euler for the hybrid format
@@
  $U^{0}_{i}\leftarrow\text{cross}\,(u_{0}(x_{i},\xi_{1},\dots,\xi_{m}))$ {Evaluate $u_{0}$ for the midpoint rule}
@@
$\bar{U}^{0}\leftarrow\text{Average}\,(U^{0})$
$\Delta t\leftarrow\alpha\dfrac{\Delta x}{\max|f^{\prime}(\bar{U}^{0})|}$.
@@
   $U^{n\pm}_{i\pm 1/2}\leftarrow\text{Reconstruction}\,(\bar{U}_{i}^{n})$ {Reconstruct values at interfaces, see Algorithm 2 }
   $a\leftarrow\text{cross}\,(\max(|f^{\prime}(U_{i+1/2}^{+})|,|f^{\prime}(U_{i+1/2}^{-})|))$ {Compute the numerical velocity}
   $H_{i\pm 1/2}\leftarrow\dfrac{f(U_{i\pm 1/2}^{+})+f(U_{i\pm 1/2}^{-})}{2\Delta x}-\dfrac{a(U_{i\pm 1/2}^{+}-U_{i\pm 1/2}^{-})}{2\Delta x}$ {Compute the left and right fluxes}
   $\bar{U}_{i}^{n+1}\leftarrow\bar{U}_{i}^{n}-\frac{\Delta t}{\Delta x}(H_{i+1/2}-H_{i-1/2})$ {Update the solution}
@@
Algorithm 2 Reconstruction
@@
  $(U_{x})_{i}^{n}\leftarrow\text{cross}\,(\text{minmod}(U_{i}-U_{i-1},U_{i+1}-U_{i}))$
@@
return $U^{n+}_{i+1/2},U^{n-}_{i+1/2}$
Algorithm 3 MUSCL scheme in the full-TT format
$U^{0}\leftarrow\text{cross}\,(u_{0}(x_{i},\xi_{1},\dots,\xi_{m}))$ {Evaluate $u_{0}$ for the midpoint rule}
$\bar{U}^{0}\leftarrow\text{Average}\,(U^{0})$
$\Delta t\leftarrow\alpha\dfrac{\Delta x}{\max|f^{\prime}(\bar{U}^{0})|}$.
for all $k\in 0,\dots,N_{t}$ do
  $U^{n\pm}\leftarrow\text{ReconstructionFTT}\,(\bar{U}^{n})$ {Reconstruct values at interfaces, see Algorithm 4 }
  $H_{L},H_{R}\leftarrow\text{NumFluxFTT}\,(U^{\pm})$ {Compute the left and right fluxes, see Algorithm 5}
  $\bar{U}^{n+1}\leftarrow\bar{U}_{i}^{n}-\frac{\Delta t}{\Delta x}(H_{R}-H_{L})$ {Update the solution}
end for
Algorithm 4 Reconstruction in full-TT format
$G_{0},G_{1},\dots,G_{m}\leftarrow\text{cores}\,(\bar{U}^{n})$ {Collect the $T$-cores of the solution}
$\tilde{G}_{0}^{L},\ \tilde{G}_{0}^{R}\leftarrow({\bf I}-{\bf L})G_{0},\ ({\bf R}-{\bf I})G_{0}$
$U^{n}_{L},\ U^{n}_{R}\leftarrow\text{TensorTrain}\,([\tilde{G}_{0}^{L},G_{1},\dots,G_{m}]),\ \text{TensorTrain}\,([\tilde{G}_{0}^{R},G_{1},\dots,G_{m}])$ {Create left- and right-shifted solution from the modified cores}
$(U_{x})^{n}\leftarrow\text{cross}\,(\text{minmod}(U_{L}^{n},U_{R}^{n}))$
$U^{n\pm}\leftarrow\bar{U}^{n}\pm\frac{1}{2}(U_{x})^{n}$
return $U^{n+},U^{n-}$
Algorithm 5 Flux computation in full-TT format
$a\leftarrow\text{cross}\,(\max(|U^{n+}|,|U^{n-}|))$ {Compute the numerical velocity}
$H\leftarrow\frac{f(U^{n+})+U^{n-}}{2}-a\frac{U^{n+}-U^{n-}}{2}$
$G_{0},G_{1},\dots,G_{m}\leftarrow\text{cores}\,(H)$ {Collect the $T$-cores of the flux}
$\tilde{G}_{0}^{L}\leftarrow{\bf L}G_{0}$ {Shift the spatial core to one index to the left}
$H_{L}\leftarrow\text{TensorTrain}\,([\tilde{G}_{0}^{L},G_{1},\dots,G_{m}])$ {Create left-shifted flux from the modified cores}
return $H_{L},H$
@@ Appendix B Expectation and variance for the exact solution
$\hat{\bm{\xi}}=(\xi_{1},\dots,\xi_{m-1}),\quad\hat{\bf v}_{L}=(v_{L,1},\dots,v_{L,m-1}),$ (51)
$\hat{\bf v}_{R}=(v_{R,1},\dots,v_{R,m-1}),\quad\hat{\bf v}=(\hat{\bf v}_{L}+\hat{\bf v}_{R})/2.$ (52)
@@
[…]
@@
$\mathbb{E}[u]=\int_{[0,1]^{m-1}}\int_{\xi_{m}=0}^{a(\hat{\bm{\xi}})}u_{R}({\bm{\xi}})p_{1}(\xi_{1})\dots p_{m}(\xi_{m})\;\text{d}\xi_{m}\;\text{d}\hat{\bm{\xi}}\\ +\int_{[0,1]^{m-1}}\int_{\xi_{m}=a(\hat{\bm{\xi}})}^{1}u_{L}({\bm{\xi}})p_{1}(\xi_{1})\dots p_{m}(\xi_{m})\;\text{d}\xi_{m}\;\text{d}\hat{\bm{\xi}}$. (54)
@@
Appendix C Shu-Osher problem: maximal ranks at various timesteps
Figures 1619 show the maximal rank in each spatial cell for the solution $\rho,u,p$ of the Shu-Osher problem at times $t\in\{0.0,0.04,0.07,0.1\}$. The results were obtained with $N_{x}=1600,N_{\xi}=20,{r_{\text{max}}}=5$ and $\varepsilon=0.01$.
Figure 16: Maximal rank in each spatial cell for $\rho,u,p$ at initial time.
Figure 17: Maximal rank in each spatial cell for $\rho,u,p$ at time $t=0.04$.
Figure 18: Maximal rank in each spatial cell for $\rho,u,p$ at time $t=0.07$.
Figure 19: Maximal rank in each spatial cell for $\rho,u,p$ at time $t=0.1$.
Acknowledgments
We would like to thank Adrian Kolb for providing the Monte-Carlo results, and the anonymous reviewer for his useful comments.